As sales manager, Joe Batista was given the following static budget report for selling expenses in the Clothing Department of Soria Company for the month of October.

SORIA COMPANY
Clothing Department
Budget Report
For the Month Ended October 31, 2020

Difference


Budget


Actual

Favorable
Unfavorable

Neither Favorable
nor Unfavorable

Sales in units

8,300

11,000

2,700

Favorable
Variable expenses
    Sales commissions

$1,992

$2,860

$868

Unfavorable
    Advertising expense

830

1,210

380

Unfavorable
    Travel expense

3,320

3,850

530

Unfavorable
    Free samples given out

1,992

1,430

562

Favorable
       Total variable

8,134

9,350

1,216

Unfavorable
Fixed expenses
     Rent

1,100

1,100

–0–

Neither Favorable nor Unfavorable
     Sales salaries

1,200

1,200

–0–

Neither Favorable nor Unfavorable
     Office salaries

900

900

–0–

Neither Favorable nor Unfavorable
     Depreciation—autos (sales staff)

400

400

–0–

Neither Favorable nor Unfavorable
       Total fixed

3,600

3,600

–0–

Neither Favorable nor Unfavorable
Total expenses

$11,734

$12,950

$1,216

Unfavorable

As a result of this budget report, Joe was called into the president’s office and congratulated on his fine sales performance. He was reprimanded, however, for allowing his costs to get out of control. Joe knew something was wrong with the performance report that he had been given. However, he was not sure what to do, and comes to you for advice.
Prepare a budget report based on flexible budget data to help Joe. (List variable costs before fixed costs.)
Answer #1
Budget Actual Difference
Sales in units 11000 11000
Variable expenses
Sales commissions 2640 2860 220 Unfavorable
Advertising expense 1100 1210 110 Unfavorable
Travel expense 4400 3850 550 Favorable
Free samples 2640 1430 1210 Favorable
Total variable 10780 9350 1430 Favorable
Fixed expenses
Rent 1100 1100 0 Neither Favorable nor Unfavorable
Sales salaries 1200 1200 0 Neither Favorable nor Unfavorable
Office salaries 900 900 0 Neither Favorable nor Unfavorable
Depreciation—autos 400 400 0 Neither Favorable nor Unfavorable
Total fixed 3600 3600 0 Neither Favorable nor Unfavorable
Total expenses 14380 12950 1430 Favorable
Workings:
Budget
Variable expenses:
Sales commissions 2640 =1992/8300*11000
Advertising expense 1100 =830/8300*11000
Travel expense 4400 =3320/8300*11000
Free samples 2640 =1992/8300*11000